### 23 Mar 2026 v1

# The Spatial Hydrodynamic Attractor: Resurgence of the Gradient Expansion

Abstract
Far-from-equilibrium kinetic systems collapse onto a hydrodynamic attractor, traditionally approximated by a gradient expansion. While temporal gradient series are non-Borel summable and require transseries completions, the analytic structure of the spatial expansion has remained elusive. Here, we derive exact closed-form Chapman–Enskog coefficients at all orders via Lagrange inversion and prove that the non-relativistic spatial gradient series, though factorially divergent, is strictly Borel summable. Furthermore, we show that this divergence originates from unbounded Galilean velocities; enforcing relativistic causality yields a convergent spatial hydrodynamic expansion with finite radius.
Introduction.—
The classical treatment of non-equilibrium systems in kinetic theory relies on a systematic gradient expansion of the distribution function, truncated at finite order to yield macroscopic fluid dynamics. The validity of this expansion is controlled by the Knudsen number, the ratio of microscopic kinetic scales to the macroscopic scales of the coarse-grained hydrodynamic description. In relativistic systems, gradients of macroscopic fields normalized by the local temperature play an analogous role. However, it is well established that the resulting gradient expansion is purely asymptotic and diverges factorially [5].
Despite this formal divergence, numerical [6] and experimental [1] evidence has shown that far-from-equilibrium systems can still be quantitatively described by hydrodynamics, suggesting the existence of a hydrodynamic attractor. Therefore, the divergent nature of the gradient expansion does not rule out the existence of such a stable manifold.
This phenomenon was first explored for temporal gradient expansions by Heller and Spaliński [10]. They demonstrated that for the highly symmetric, longitudinally expanding quark-gluon plasma (Bjorken flow) within Müller-Israel-Stewart framework, the fluid dynamics reduce to a set of ordinary differential equations. They showed that despite the factorial divergence of the hydrodynamic series at large orders, the attractor rigorously exists. Its exact reconstruction, requires generalized Borel resummation; the series is non-Borel summable along the positive real axis due to singularities corresponding to the decay of non-hydrodynamic modes, necessitating a full transseries completion.
This decaying of non-hydrodynamic modes (attractor paradigm) was subsequently generalized by Romatschke [14], who demonstrated that conformal Bjorken flow across three distinct microscopic theories: a variant of second-order BRSSS hydrodynamics [3], Boltzmann theory, and strongly coupled $\mathcal{N}=4$ supersymmetric Yang-Mills (via AdS/CFT), universally collapses onto a hydrodynamic attractor.
These foundational studies focused exclusively on the temporal (longitudinal) gradient expansion, establishing that its divergence universally requires generalized Borel resummation. Unlike temporal evolution, spatial gradients trigger short-wavelength ultraviolet pathologies, such as the Burnett instability; the Borel summability and resurgent structure of the spatial gradient series has not been systematically studied.
In this letter, we show that for non-relativistic Boltzmann kinetic theory, the spatial gradient expansion is factorially divergent but strictly Borel summable, unambiguously reconstructing the spatial attractor; moreover, once relativistic causality is imposed on the velocity space, the divergence is cured entirely, yielding a strictly convergent series with a finite radius.
Spatial Attractor of the Kinetic Equation.—
Starting from the one-dimensional kinetic Bhatnagar-Gross-Krook (BGK) equation [4] and setting the particle mass $m$, thermal energy $k_{B}T$, and relaxation time $\tau$ to unity, we write the non-dimensionalized kinetic equation for the distribution function $f(x,v,t)$ as
$\partial_{t}f=-v\,\partial_{x}f-(f-f^{\rm eq}),$ (1)
where the local equilibrium is the Maxwellian $f^{\rm eq}=n(x,t)\,(2\pi)^{-1/2}\,e^{-v^{2}/2}$, and $n(x,t)=\int_{-\infty}^{\infty}f\,dv$ is the locally conserved particle density. Using the velocity moments of the distribution function,
$M_{l}(x,t)=\int_{-\infty}^{\infty}v^{l}f\,dv,\quad l=0,1,2,\ldots,$
in (1) yields an infinite, unclosed hierarchy of moment equations. The equilibrium moments take the Gaussian form,
$M_{l}^{\rm eq}(x,t)=\begin{cases}n(x,t)(l-1)!!&l\text{ even},\\ 0&l\text{ odd}\end{cases}.$
Following the machinery typically used in the method of invariant manifolds [8], the infinite hierarchy is compactly encoded in the generating function $Z(\lambda,x,t)=\int_{-\infty}^{\infty}e^{-i\lambda v}f\,dv$. Applying the spatial Fourier transform to the even (real) and odd (imaginary) parts of $Z$ parametrizes the hydrodynamic manifold so that all moments depend on the Fourier-transformed conserved density $\hat{n}$,
$\partial_{t}\hat{Z}_{\text{re}}$ $=ik\,\partial_{\lambda}\hat{Z}_{\text{im}}-\hat{Z}_{\text{re}}+\hat{n}\,e^{-\lambda^{2}/2},$ (2)
$\partial_{t}\hat{Z}_{\text{im}}$ $=-ik\,\partial_{\lambda}\hat{Z}_{\text{re}}-\hat{Z}_{\text{im}}.$ (3)
In Fourier space this means $\hat{Z}_{\text{re}}=\hat{\Theta}_{\text{re}}(\lambda,k^{2})\,\hat{n}$ and $\hat{Z}_{\text{im}}=ik\,\hat{\Theta}_{\text{im}}(\lambda,k^{2})\,\hat{n}$, where the factor $ik$ in $\hat{Z}_{\text{im}}$ reflects the odd parity of the flux. The central requirement is the dynamic invariance condition: the macroscopic and microscopic time derivatives must agree exactly on this manifold [9, 11], bridging the kinetic and hydrodynamic descriptions in the spirit of Hilbert’s sixth problem.
We now briefly outline the approach developed in Ref. [12]. Imposing the invariance condition yields a system of two first-order ODEs in $\lambda$ for $\hat{\Theta}_{\text{re,im}}$. Eliminating $\hat{\Theta}_{\text{re}}$ and introducing the frequency function $\hat{\Omega}(\lambda,k^{2})$ via $\partial_{\lambda}\hat{\Theta}_{\text{im}}=-k^{-2}\,\hat{\Omega}(\lambda,k^{2})\,e^{-\lambda^{2}/2}$, one obtains a single exact ODE for $\hat{\Omega}$:
[…]
subject to the initial conditions
$\hat{\Omega}(0,k^{2})=\hat{\omega}(k),\qquad\left.\partial_{\lambda}\hat{\Omega}\right|_{\lambda=0}=0.$ (5)
Equation (4) is the exact, non-perturbative hydrodynamic equation for this system; however, the initial conditions (5) do not fix $\hat{\omega}(k)$ uniquely, and in prior work the physical branch was isolated numerically by an iterative “pullout” procedure [12]. Here, we show that this phase-space dimensionality reduction can be carried out entirely algebraically. Since the frequency function is even and analytic in $\lambda$, we can expand it in a power series
$\hat{\Omega}(\lambda,k^{2})=\hat{\omega}+\sum_{n=1}^{\infty}\frac{\lambda^{2n}\,\hat{\omega}_{2n}(\hat{\omega},k^{2})}{(2n)!\,k^{2n}},$ (6)
Substituting Eq. (6) into Eq. (4), one finds that the coefficients factorize as $\hat{\omega}_{2n}=(\hat{\omega}+1)P_{n}(\hat{\omega},k^{2})$, where $P_{n}$ is a family of spectral polynomials satisfying the three-term recurrence ($n\geq 2$):
$P_{n}=\bigl[(\hat{\omega}+1)^{2}+(4n{-}3)\,k^{2}\bigr]\,P_{n-1}\\ -k^{4}\,(2n{-}2)(2n{-}3)\,P_{n-2},$
with $P_{0}=1$ and $P_{1}=\hat{\omega}(\hat{\omega}+1)+k^{2}$. The three-term structure is a direct consequence of the Hermite-type differential operator $\partial_{\lambda}^{2}-2\lambda\partial_{\lambda}$ in Eq. (4). The root $\hat{\omega}=-1$ is the kinetic eigenvalue of the BGK equation (1), which trivially defines an invariant manifold at every truncation order. Truncating the hierarchy at order $n$ by setting $\hat{\omega}_{2n}=0$ yields the algebraic equation $P_{n}(\hat{\omega},k^{2})=0$. Here, we explore this polynomial systematically via numerical continuation using Auto-07p [7] to isolate the physical root $\hat{\omega}(k)$ that defines the hydrodynamic attractor. As shown in Fig. 2, the spectral polynomial branches $P_{n}=0$ converge to the non-perturbative attractor with increasing truncation order; each branch terminates at a fold bifurcation beyond a critical wavenumber $k_{c}(n)$ that grows with $n$.
Exact Chapman-Enskog Coefficients.—
The differential operator $\mathcal{L}=\partial_{\lambda}^{2}-2\lambda\,\partial_{\lambda}$ on the right-hand side of Eq. (4) satisfies the conjugation identity $\mathcal{L}[e^{\lambda^{2}/2}\Phi]=e^{\lambda^{2}/2}[\Phi^{\prime\prime}+(1-\lambda^{2})\Phi]$. Since the factor $(1-\lambda^{2})$ matches the potential multiplying $(\hat{\Omega}+1)$, we introduce the substitution
$\hat{\Omega}(\lambda)=(\hat{\omega}+1)\,e^{\lambda^{2}/2}\,\Phi(\lambda)-1$
which eliminates all $\lambda$-dependent coefficients, reducing Eq. (4) to the constant-coefficient ODE
$\Phi^{\prime\prime}(\lambda)-A\Phi(\lambda)=-\frac{A}{\hat{\omega}+1}e^{-\lambda^{2}/2},\quad A=\frac{(\hat{\omega}+1)^{2}}{k^{2}},$ (7)
with $\Phi(0)=1$ and $\Phi^{\prime}(0)=0$.
The homogeneous solutions of Eq. (7) are $e^{\pm\sqrt{A}\,\lambda}$. Since $f(x,v,t)$ is integrable in $v$, the Riemann–Lebesgue lemma requires $Z(\lambda)$ to remain bounded as $|\lambda|\to\infty$; $\hat{\Omega}$ may grow at most as $e^{\lambda^{2}/2}$, and consequently $\Phi(\lambda)$ must be bounded. This rules out both homogeneous solutions.
The unique bounded particular solution is obtained by convolving the source with the free-space Green’s function $G(\lambda,s)=-\frac{1}{2\sqrt{A}}\,e^{-\sqrt{A}\,|\lambda-s|}$ of the operator $d^{2}/d\lambda^{2}-A$:
$\Phi(\lambda)=\frac{\sqrt{A}}{2(\hat{\omega}+1)}\int_{-\infty}^{\infty}e^{-\sqrt{A}\,|\lambda-s|}\;e^{-s^{2}/2}\,ds.$ (8)
Evaluating Eq. (8) at $\lambda=0$ and imposing $\Phi(0)=1$, the two-sided exponential kernel is recast as a velocity-space resolvent via its Fourier representation $\tfrac{\sqrt{A}}{2}\,e^{-\sqrt{A}|s|}=\frac{1}{2\pi}\!\int\!\frac{A}{A+v^{2}}\,e^{ivs}\,dv$, followed by Gaussian $s$-integration. This yields the exact self-consistency condition
$\hat{\omega}+1=\int_{-\infty}^{\infty}\frac{A}{A+v^{2}}\;\frac{e^{-v^{2}/2}}{\sqrt{2\pi}}\,dv,$
For small $k$ (equivalently large $A$), the resolvent is expanded as a geometric series, $\frac{A}{A+v^{2}}=\sum_{m=0}^{\infty}(-1)^{m}(v^{2}/A)^{m}$, and integrated term by term using the Gaussian moments $\langle v^{2m}\rangle=(2m{-}1)!!$ :
$\hat{\omega}$ $=\sum_{m=1}^{\infty}(-1)^{m}\,(2m{-}1)!!\;\frac{k^{2m}}{(1{+}\hat{\omega})^{2m}}$
$=\sum_{m=1}^{\infty}(-1)^{m}\,(2m{-}1)!!\;x^{m}\;:=\;F(x),$ (9)
where $x=k^{2}(1{+}\hat{\omega})^{-2}=A^{-1}$. Since $x$ itself depends on $\hat{\omega}$, Eq. (Exact Chapman-Enskog Coefficients.—) is an implicit relation for $\hat{\omega}(k)$. The Chapman–Enskog (CE) coefficients $\hat{\omega}=\sum_{n=1}^{\infty}a_{2n}\,k^{2n}$ are extracted by Lagrange inversion:
$a_{2n}=\frac{1}{n}\,[x^{n-1}]\bigl\{F^{\prime}(x)\,(1{+}F(x))^{-2n}\bigr\},$ (10)
where $x^{m}]\,g(x)$ denotes the coefficient of $x^{m}$ in the power series of $g$ (see Chap. 5 of Ref. [15] for details on coefficient extraction notation and Lagrange inversion of power series).
Figure 1: Absolute values of the exact CE coefficients $|a_{2n}|$ from Eq. (10) (black circles, left axis) and ratios $r_{n}=|a_{2(n+1)}/a_{2n}|$ (blue squares, right axis). The convex shape of $|a_{2n}|$ on a log scale signals factorial divergence, while the asymptotically linear growth $r_{n}\sim 2n$ implies zero radius of convergence. This factorial scaling is the spatial-gradient analog of the temporal divergence figure in Ref. [10]; crucially, the spatial series is strictly Borel summable.
Equation (10) provides, to our knowledge, novel closed-form expressions for the exact CE transport coefficients at all orders, determined entirely by the equilibrium velocity moments (2m1)!!(2m{-}1)!!. The explicit evaluation of the first few coefficients yields $a_{2}=-1,a_{4}=1,a_{6}=-4,a_{8}=27,a_{10}=-248,a_{12}=2830,a_{14}=-38232$, matching the sequence reported in Ref. [12]. As shown in Fig. 1, the coefficients grow factorially and the ratio $r_{n}=|a_{2(n+1)}/a_{2n}|$ increases roughly linearly; we later show that $r_{n}\sim 2n$, implying zero radius of convergence for the spatial CE series.
Resurgence and Borel Summability.—
The factorial divergence of the CE coefficients originates from the nature of the series $F(x)=\sum_{m=1}^{\infty}(-1)^{m}(2m{-}1)!!\,x^{m}$. The classical approach to analyzing $F$ is through its Borel transform, $\mathcal{B}[F]$. Using the identities $(2m{-}1)!!/m!=\binom{2m}{m}/2^{m}$ and $\sum_{m=0}^{\infty}\binom{2m}{m}y^{m}=(1-4y)^{-1/2}$, we obtain
$\mathcal{B}[F](\sigma)=\sum_{m=1}^{\infty}\frac{(-1)^{m}(2m{-}1)!!}{m!}\,\sigma^{m}=\frac{1}{\sqrt{1+2\sigma}}-1.$
To understand the analytic structure of $\mathcal{B}[F]$, we note that $F$ is recovered from its Borel transform via the directional Laplace transform: $F(x)=\int_{0}^{+\infty}e^{-\sigma/x}\mathcal{B}[F](\sigma)\,d\sigma$. Since the Borel transform has a unique singularity on the negative real axis at $\sigma_{\star}=-1/2$, the Laplace contour along the positive real axis is entirely unobstructed. Consequently, the CE series is strictly Borel summable (exhibiting no Stokes phenomenon), and the Borel sum unambiguously reconstructs the exact dispersion relation $\hat{\omega}(k)$ without the need for lateral contour deformation or transseries completion.
Furthermore, the large-order growth of the CE coefficients follows the asymptotic behavior $a_{2n}\sim(-1)^{n}\,n!\,2^{n}$, rigorously explaining the ratio $r_{n}\sim 2n$ observed in Fig. 1.
This strict summability stands in contrast to the temporal gradient expansion of Bjorken flow. In Ref. [10] the Borel transform of the Müller-Israel-Stewart temporal gradient series possesses its leading singularity at $\xi_{0}=3/(2C_{\tau\Pi})>0$ on the positive real axis, corresponding to the decay rate of the non-hydrodynamic mode. In that case, the Laplace contour is obstructed, rendering the series non-Borel summable and requiring a full transseries completion.
In the spatial case, the divergence originates from an ultraviolet microscopic effect, namely the unbounded velocity tail of the equilibrium distribution. This is a fundamentally different mechanism from the temporal case, where the divergence is due to a macroscopic non-hydrodynamic mode. This difference explains both the negative sign, as the alternating CE coefficients reflect the competition between even and odd velocity moments, and the strict Borel summability. The Burnett instability is therefore not a fundamental pathology but a truncation artifact of this alternating divergent series, fully resolved by Borel resummation.
Figure 2: Hydrodynamic dispersion relation $\hat{\omega}(k)$. Solid colored curves show the spectral polynomial branches $P_{n}(\hat{\omega},k^{2})=0$ for $n=1,2,20,50$, converging to the non-perturbative attractor with increasing truncation order; faded segments beyond the critical wavenumber $k_{c}$ (filled circles) indicate the unphysical region past the fold bifurcation, with $k_{c}\approx 0.47,0.58,0.94,1.03$ for $n=1,2,20,50$ respectively. The diagonal Padé approximant of order $[14/14]$, constructed from the first 30 Borel-transformed CE coefficients (red), exactly reconstructs the attractor. Finite-order CE truncations CE(2) (dashed) and CE(4) (dash-dotted) are both divergent.
Figure 2 confirms this: the Borel–Padé resummation of the CE series exactly reconstructs the non-perturbative attractor, while the spectral polynomial branches $P_{n}=0$ converge slowly to the hydrodynamic manifold with increasing truncation order. By contrast, the first-order CE truncation (classical diffusion), $\mathrm{CE^{2}}:\hat{\omega}=-k^{2}$, and the second-order Burnett-type approximation, $\mathrm{CE^{4}}:\hat{\omega}=-k^{2}+k^{4}$, diverge at moderate $k$.
Relativistic BGK.—
Since the factorial divergence originates from the unbounded velocity tail, it is tempting to consider a relativistic kinetic equation, imposing relativistic causality, resulting in the velocity phase space being bounded by the speed of light.
A simple case of the relativistic BGK is considered: the Anderson–Witting model [2] with the relaxation rate ${p^{\mu}u_{\mu}}/{\tau}$ reads,
$p^{\mu}\partial_{\mu}f=-\frac{p^{\mu}u_{\mu}}{\tau}(f-f^{\rm eq}),$ (11)
Here, we consider the fluid rest frame, $u^{\mu}=(1,0,0,0)$, $p^{0}=E=\sqrt{p^{2}+m_{0}^{2}}$ is the particle energy, $p^{1}=p$ is the spatial momentum, and $m_{0}$ is the rest mass. We set the relaxation time $\tau=1$ and use the streaming particle velocity $v=p/E$ to recast the Eq. (11) in the form $\partial_{t}f+v\,\partial_{x}f=-(f-f^{\rm eq})$. This form is mathematically identical to the non-relativistic BGK equation (1). However, the velocity phase space is now bounded by the speed of light, $v\in[-1,1]$.
The velocity moments are now also bounded as
$M_{l}(x,t)=\int_{-1}^{1}v^{l}f(x,v,t)\,dv,\qquad l=0,1,2,\ldots$
The equilibrium distribution is $f^{\rm eq}=n(x,t)W(v)$, where $W(v)$ is any normalized, symmetric weight function on [1,1][-1,1] (e.g., the Maxwell–Jüttner distribution), provides the equilibrium moments $M_{l}^{\rm eq}(x,t)=n(x,t)\,\mu_{l}$, where
$\mu_{l}=\int_{-1}^{1}v^{l}W(v)\,dv.$
By symmetry $\mu_{2m+1}=0$ and normalization fixes $\mu_{0}=1$.
The evolution of the generating function $\hat{Z}$ takes a similar form to Eqs. (2)–(3) with one change: the $e^{-\lambda^{2}/2}$ term is now $\hat{W}(\lambda)=\int_{-1}^{1}e^{-i\lambda v}W(v)\,dv$. Imposing the invariance condition on the Fourier-transformed kinetic equation, solving for $\hat{f}$, and integrating over $v\in[-1,1]$ yields the self-consistency condition
$1=\int_{-1}^{1}\frac{W(v)}{1+\hat{\omega}+ikv}\,dv.$
[…]
$F_{\rm rel}(x)=\sum_{m=1}^{\infty}(-1)^{m}\mu_{2m}\,x^{m}.$ (12)
Equation (12) is the relativistic analog of Eq. (Exact Chapman-Enskog Coefficients.—), and CE coefficients follow from the same Lagrange inversion (10) replacing $F$ with $F_{\rm rel}$. The crucial difference is that the bounded support $v\in[-1,1]$ guarantees $\mu_{2m}\leq 1$ for all $m$, in contrast to the factorially growing Gaussian moments (2m1)!!(2m{-}1)!!. Consequently, $F_{\rm rel}(x)$ has a strictly non-zero radius of convergence, and by the inverse function theorem the CE series $\hat{\omega}(k)=\sum a_{2n}k^{2n}$ also converges. The spatial gradient expansion is therefore strictly convergent: relativistic causality eliminates the factorial divergence at its source.
Conclusions and Outlook.—
We have shown that for non-relativistic BGK kinetic theory, the spatial gradient expansion of the hydrodynamic series diverges factorially; however, it is strictly Borel summable. Recursive spectral polynomials as well as the Borel–Padé resummation reconstruct the non-perturbative hydrodynamic attractor, while finite-order CE truncations diverge at moderate wavenumbers. Previous works [10, 14] have shown that the temporal gradient expansion is also factorially divergent, but non-Borel summable, requiring a full transseries completion to reconstruct the attractor.
In addition, we derived closed-form expressions for the exact CE transport coefficients at all orders via Lagrange inversion, and showed that imposing relativistic causality on the velocity phase space cures the divergence entirely, yielding a convergent series with finite radius.
Taken together with the results of Refs. [10] and [14], our findings suggest a unifying picture: the hydrodynamic gradient expansion, whether temporal or spatial, is always Borel summable, and its resummation invariably reconstructs the unique non-perturbative attractor defined by the (slow) invariant manifold.
Our results suggest that the rigorous passage from kinetics to hydrodynamics, a central aspect of Hilbert’s sixth problem [9], need not rely on a convergent perturbative expansion. Instead, we conjecture that hydrodynamics can be systematically derived from kinetic theory through Borel resummation of the factorially divergent gradient series, providing a non-perturbative route to the hydrodynamic limit.
Several natural extensions include: (i) extending beyond the BGK operator to the full Boltzmann collision integral with nonlinear collision kernels, and (ii) repeating the relativistic analysis with a corrected relaxation-time approximation [13] that respects microscopic conservation laws.